# Structure-Aware Correspondence Learning for Relative Pose Estimation

### v1

@@ Abstract
Relative pose estimation provides a promising way for achieving object-agnostic pose estimation. Despite the success of existing 3D correspondence-based methods, the reliance on explicit feature matching suffers from small overlaps in visible regions and unreliable feature estimation for invisible regions. Inspired by humans’ ability to assemble two object parts that have small or no overlapping regions by considering object structure, we propose a novel Structure-Aware Correspondence Learning method for Relative Pose Estimation, which consists of two key modules. First, a structure-aware keypoint extraction module is designed to locate a set of kepoints that can represent the structure of objects with different shapes and appearance, under the guidance of a keypoint based image reconstruction loss. Second, a structure-aware correspondence estimation module is designed to model the intra-image and inter-image relationships between keypoints to extract structure-aware features for correspondence estimation. By jointly leveraging these two modules, the proposed method can naturally estimate 3D-3D correspondences for unseen objects without explicit feature matching for precise relative pose estimation. Experimental results on the CO3D, Objaverse and LineMOD datasets demonstrate that the proposed method significantly outperforms prior methods, i.e., with $5.7^{\circ}$ reduction in mean angular error on the CO3D dataset.
@@ 1 Introduction
Object pose estimation aims to estimate the 3D translation and 3D rotation of an object from a single image. It plays a crucial role in many real-world applications such as augmented reality (AR) [2, 26], robotic manipulation [27, 45, 44] and autonomous driving [4, 9], drawing increasing attention in recent years. Early works [46, 41, 12, 13] mainly focus on instance-level pose estimation, where the model is trained to estimate the pose for a specific object. However, these methods cannot generalize to other objects. Consequently, the category-level pose estimation [42] is introduced, where the model is trained to estimate the pose for different instances within the same category. Nevertheless, they can hardly generalize to unseen objects of other categories, limiting their application potentials.
@@
To further improve the generalization ability, recent works [48, 19, 49] have shifted towards relative pose estimation for unseen objects. It requires only a single reference image of a novel object to estimate the poses of new images, as shown in Figure 1 (a), providing a promising way to achieve object-agnostic pose estimation. Existing relative pose estimation methods can be broadly categorized into three types, 2D correspondence-based methods [36, 33], hypothesis-and-verification-based methods [48, 19, 49] and 3D correspondence-based methods [49, 50]. 2D correspondence-based methods [36, 33] extract keypoints directly from two images and compute the relative pose based on keypoint matching. However, as shown in Figure 1 (b), these methods suffer from small overlapping regions caused by large pose variations, making it difficult to establish reliable correspondence. Hypothesis-and-verification-based methods [48, 19, 49] sample a large number of pose hypotheses and then evaluate the score of each hypothesis through global feature matching or score regression. Nevertheless, these methods rely on a discrete sampling process and fail to adequately model the continuous pose space, which can only account for coarse pose estimation. Moreover, the verification of numerous pose hypotheses incurs significant computational costs. 3D correspondence-based methods [49, 50] lift 2D features into 3D voxel features and then conduct 3D-3D matching to estimate the pose, as shown in Figure 1 (c). Notably, these methods can establish correspondences even in invisible regions, showing promising potential. However, it’s difficult to infer 3D features of invisible regions from 2D surface features without extra information, resulting in unreliable 3D matching for invisible regions. Moreover, dense 3D-3D matching process incurs high computational costs due to the cubic complexity.
@@
  • Experimental results on three challenging datasets, CO3D [30], Objaverse [6] and LineMOD [14], demonstrate the state-of-the-art performance of our method.
@@ 2.1 Instance-Level Object Pose Estimation
Instance-level pose estimation methods [46, 41, 12, 13] predict the 6D pose of specific known objects by leveraging CAD models. Recent approaches include correspondence, template, voting, and direct regression methods. Correspondence-based methods [29, 18, 47] establish matches between inputs and CAD models, then estimate pose via PnP [7] or similar algorithms. Template-based methods [37, 17] match inputs to pre-defined templates, treating pose estimation as a matching problem. Voting-based methods [24, 39] aggregate votes from pixels or points, either by keypoints or direct pose prediction. Regression-based methods [8, 22] directly predict 6D poses from images or depth data, simplifying the pipeline. Although effective for known objects, these methods struggle to generalize to unseen objects.
@@ 2.2 Category-level Object Pose Estimation
To extend pose estimation beyond specific instances, researchers have developed category-level methods [42, 38, 20, 21, 23, 31, 32] for estimating object poses within predefined categories without CAD models. These methods fall into two categories: shape prior-based and shape prior-free. Shape prior-based approaches [42, 38, 20] utilize CAD-derived priors for alignment or direct regression of poses. In contrast, shape prior-free methods [21, 23, 5, 31, 32] remove reliance on priors, learning features directly from input data to estimate poses. Although category-level methods improve generalization over instance-specific methods, they still struggle to generalize to unseen categories.
@@ 2.3 Relative Object Pose Estimation
To enhance the generalization of pose estimation, recent work [48, 49, 19, 50] has focused on relative pose estimation for unseen objects. Unlike instance-specific methods, these approaches require only a single reference image of a novel object to estimate the relative pose of image, making them highly suitable for applications where data acquisition is costly. Current methods can be broadly categorized into three types: 2D correspondence-based methods [36, 33, 25], hypothesis-and-verification-based methods [48, 19, 49, 28], and 3D correspondence-based methods [49, 50]. 2D correspondence-based methods [36, 33, 25] establish correspondences between keypoints from both images and use these to compute relative pose. Learned feature-based approaches like SuperGlue [33] and LoFTR [36] have achieved robustness in feature matching under moderate viewpoint changes and lighting variations. However, with only a single reference image and substantial viewpoint differences, these methods struggle due to their sensitivity, significantly affecting accuracy [48, 49]. Hypothesis-and-verification-based methods [48, 19, 49, 28] mitigate these challenges by generating multiple pose hypotheses over the rotation space and evaluating them through similarity networks, as in RelPose [48] and RelPose++ [19]. While effective in handling larger viewpoint differences, these methods require extensive sampling and verification, resulting in high computational costs that restrict their suitability for real-time applications. As an alternative, 3D correspondence-based methods [50, 49] lift 2D features into 3D voxel features and then conduct 3D-3D matching to estimate the pose. For example, DVMNet [50] uses lifted 3D voxel features to facilitate matching even in invisible regions. However, it’s difficult to infer 3D features solely from 2D surface image features, which leads to unreliable 3D matching. Moreover, the dense 3D-3D matching incurs high computational costs due to the cubic complexity.
@@ 3.1 Overview
We tackle the problem of estimating the relative pose between a query image and a reference image belonging to previously unseen object categories. Specifically, we denote the set of object categories available during training as $\omega_{\text{train}}$ and the set of categories used during testing as $\omega_{\text{test}}$, where $\omega_{\text{train}}\cap\omega_{\text{test}}=\emptyset$. This setup introduces several key challenges, including significant viewpoint variations between the images, minimal overlapping regions, and the generalization to novel object categories. Since the translation component of 6D pose can be reliably estimated through existing 2D detection techniques [11, 16], we focus on addressing the more challenging 3D rotation estimation, which is the same with the mainstream methods [48, 49, 50].
The framework of our method is shown in Figure 2 (a). Given the query image $I_{q}$ and the reference image $I_{r}$, we first employ a shared feature extractor to obtain feature maps $\mathbf{F}_{q}$ and $\mathbf{F}_{r}$. These feature maps are subsequently passed through symmetric attention blocks [40] to enhance the image features mutually. Afterward, we use the proposed structure-aware keypoint extraction module to extract keypoints independently from each image. Then, by leveraging our proposed structure-aware correspondence estimation module, we facilitate both intra-image and inter-image feature interactions for the keypoints extracted from the query image, enabling structure-aware feature aggregation. With these updated keypoint features, we lift 2D keypoints to 3D space and regress their corresponding 3D coordinates within the reference coordinate system. Finally, based on the established correspondences, the relative rotation $\Delta\mathbf{R}$ is obtained via a weighted Singular Value Decomposition (wSVD) algorithm [3].
@@ 3.2 Feature Extraction
We first utilize a pre-trained backbone [43] to extract image features from both the query and reference images, yielding feature maps $\mathbf{F}_{q},\mathbf{F}_{r}\in\mathbb{R}^{H\times W\times C}$. These feature maps are then processed through a two-step attention mechanism with Multi-Head Self-Attention (MHSA) and Multi-Head Cross-Attention (MHCA) layers to update features:
@@
The same operations are performed for $\mathbf{F}_{r}$. Unlike previous works [49, 50], the parameters in the MHSA and MHCA modules are shared across the query ($q$) and reference ($r$) images, ensuring consistency between the learned features. After $L$ layers of such interactions, we obtain the final feature maps $\mathbf{F}_{q}^{(L)}$ and $\mathbf{F}_{r}^{(L)}$.
@@ 3.3 Structure-Aware Keypoint Extraction
where $\phi_{l}(\cdot)$ denotes the feature map from the $l$-th layer of the VGG network [34], and $\lambda_{1},\lambda_{2}$ are weighting factors. The $L_{2}$ loss ensures pixel accuracy, while the perceptual loss maintains semantic consistency. By reconstructing the image, we can optimize keypoint distribution end-to-end, ensuring these keypoints cover semantically rich regions of the object’s surface and enhance structural representation.
@@ 3.5 Pose Estimation via 3D Correspondences
After obtaining the 3D keypoint correspondences, following previous work [50], we employ a weighted Singular Value Decomposition (wSVD) approach to solve the relative rotation $\Delta\textbf{R}$:
@@
To solve the minimization problem for estimating the optimal rotation matrix $\Delta\textbf{R}$, we follow the approach of utilizing SVD method to derive the optimal alignment between two sets of 3D keypoints, as used in prior works on point cloud registration [1]. Specifically, we first compute the covariance matrix $\mathbf{H}$ using the query and reference 3D keypoints along with their confidence scores:
@@
where $q(\Delta\mathbf{R})$ and $q(\Delta\mathbf{R}_{\text{gt}})$ denote the 6D representation for $\Delta\mathbf{R}$ and $\Delta\mathbf{R}_{\text{gt}}$ introduced in  [51].
@@ 4 Experiment
SuperGlue [33] 2D 67.2 45.2 37.7 102.4 15.1 12.1 64.8 26.2 14.3
LoFTR [36] 2D 77.5 37.9 33.1 134.1 9.6 7.7 84.5 24.2 13.5
ZSP [10] 2D 87.5 25.7 14.6 107.2 4.2 1.5 78.6 10.7 2.7
RelPose [48] H&V 50.0 64.2 48.6 80.4 20.8 6.7 58.3 26.1 7.0
RelPose++ [19] H&V 38.5 77.0 69.8 33.5 72.3 42.9 46.6 42.5 15.1
3DAHV [49] H&V 3D 28.5 83.5 71.0 28.1 78.6 58.4 41.7 61.5 29.9
DVMNet [50] 3D 19.9 85.9 62.3 20.2 81.5 57.2 36.8 55.1 23.8
@@
Datasets. Following previous works [49, 50], we evaluate our method on CO3D [30], Objaverse [6] and LineMOD [14], which are widely-used datasets for relative pose estimation. These datasets include diverse synthetic and real data across various object categories. The CO3D dataset contains 18,619 video sequences spanning 51 categories. Following [30], we train on 41 categories and test on 10 unseen categories to evaluate generalization. The Objaverse dataset consists of synthetic images rendered from 3D models across diverse viewpoints. We select 128 objects for testing and reserve the remaining for training. For LineMOD, we use calibrated real images of 13 household objects. The test set includes 5 objects, which are excluded from training to ensure complete separation.
Implementation Details. The Adam optimizer [15] is employed with an initial learning rate of $2\times 10^{-4}$, which decays by a factor of 0.1 every 200 epochs. The model is trained for 400 epochs with a batch size of 80. All experiments are conducted on 4 NVIDIA RTX3090 GPUs, taking roughly 36 hours. Following previous works [48, 19, 49, 50], we crop the object from the image by utilizing the ground truth bounding box. Further details can be found in the supplementary material.
Evaluation Metrics. Following previous works [49, 50], we evaluate our model using two metrics: mean angular error (mAE) and accuracy under predefined thresholds. The angular error $\theta$ between the predicted rotation $\Delta R$ and the ground truth $\Delta R_{\text{gt}}$ is calculated as:

### v2

@@ Abstract
Relative pose estimation provides a promising way for achieving object-agnostic pose estimation. Despite the success of existing 3D correspondence-based methods, the reliance on explicit feature matching suffers from small overlaps in visible regions and unreliable feature estimation for invisible regions. Inspired by humans’ ability to assemble two object parts that have small or no overlapping regions by considering object structure, we propose a novel Structure-Aware Correspondence Learning method for Relative Pose Estimation, which consists of two key modules. First, a structure-aware keypoint extraction module is designed to locate a set of kepoints that can represent the structure of objects with different shapes and appearance, under the guidance of a keypoint based image reconstruction loss. Second, a structure-aware correspondence estimation module is designed to model the intra-image and inter-image relationships between keypoints to extract structure-aware features for correspondence estimation. By jointly leveraging these two modules, the proposed method can naturally estimate 3D-3D correspondences for unseen objects without explicit feature matching for precise relative pose estimation. Experimental results on the CO3D, Objaverse and LineMOD datasets demonstrate that the proposed method significantly outperforms prior methods, i.e., with $5.7^{\circ}$ reduction in mean angular error on the CO3D dataset. Our code is available at https://github.com/Cyhhzo02/SAC-Pose-code.
@@ 1 Introduction
Object pose estimation aims to estimate the 3D translation and 3D rotation of an object from a single image. It plays a crucial role in many real-world applications such as augmented reality (AR) [azuma1997survey, marchand2015pose], robotic manipulation [mousavian20196, wu2020grasp, wen2022you] and autonomous driving [chen2017multi, geiger2012we], drawing increasing attention in recent years. Early works [xiang2017posecnn, wang2019densefusion, he2020pvn3d, he2021ffb6d] mainly focus on instance-level pose estimation, where the model is trained to estimate the pose for a specific object. However, these methods cannot generalize to other objects. Consequently, the category-level pose estimation [wang2019normalized] is introduced, where the model is trained to estimate the pose for different instances within the same category. Nevertheless, they can hardly generalize to unseen objects of other categories, limiting their application potentials.
@@
To further improve the generalization ability, recent works [zhang2022relpose, lin2024relpose++, zhao20233d] have shifted towards relative pose estimation for unseen objects. It requires only a single reference image of a novel object to estimate the poses of new images, as shown in Figure 1 (a), providing a promising way to achieve object-agnostic pose estimation. Existing relative pose estimation methods can be broadly categorized into three types, 2D correspondence-based methods [sun2021loftr, sarlin2020superglue], hypothesis-and-verification-based methods [zhang2022relpose, lin2024relpose++, zhao20233d] and 3D correspondence-based methods [zhao20233d, zhao2024dvmnet]. 2D correspondence-based methods [sun2021loftr, sarlin2020superglue] extract keypoints directly from two images and compute the relative pose based on keypoint matching. However, as shown in Figure 1 (b), these methods suffer from small overlapping regions caused by large pose variations, making it difficult to establish reliable correspondence. Hypothesis-and-verification-based methods [zhang2022relpose, lin2024relpose++, zhao20233d] sample a large number of pose hypotheses and then evaluate the score of each hypothesis through global feature matching or score regression. Nevertheless, these methods rely on a discrete sampling process and fail to adequately model the continuous pose space, which can only account for coarse pose estimation. Moreover, the verification of numerous pose hypotheses incurs significant computational costs. 3D correspondence-based methods [zhao20233d, zhao2024dvmnet] lift 2D features into 3D voxel features and then conduct 3D-3D matching to estimate the pose, as shown in Figure 1 (c). Notably, these methods can establish correspondences even in invisible regions, showing promising potential. However, it’s difficult to infer 3D features of invisible regions from 2D surface features without extra information, resulting in unreliable 3D matching for invisible regions. Moreover, dense 3D-3D matching process incurs high computational costs due to the cubic complexity.
@@
  • Experimental results on three challenging datasets, CO3D [reizenstein2021common], Objaverse [deitke2023objaverse] and LineMOD [hinterstoisser2013model], demonstrate the state-of-the-art performance of our method.
@@ 2.1 Instance-Level Object Pose Estimation
Instance-level pose estimation methods [xiang2017posecnn, wang2019densefusion, he2020pvn3d, he2021ffb6d] predict the 6D pose of specific known objects by leveraging CAD models. Recent approaches include correspondence, template, voting, and direct regression methods. Correspondence-based methods [rad2017bb8, li2019cdpn, zakharov2019dpod] establish matches between inputs and CAD models, then estimate pose via PnP [fischler1981random] or similar algorithms. Template-based methods [sundermeyer2018implicit, li2022dcl] match inputs to pre-defined templates, treating pose estimation as a matching problem. Voting-based methods [liu2021kdfnet, tian2020robust] aggregate votes from pixels or points, either by keypoints or direct pose prediction. Regression-based methods [gao20206d, lin20216d] directly predict 6D poses from images or depth data, simplifying the pipeline. Although effective for known objects, these methods struggle to generalize to unseen objects.
@@ 2.2 Category-level Object Pose Estimation
To extend pose estimation beyond specific instances, researchers have developed category-level methods [wang2019normalized, tian2020shape, lin2022category, lin2023vi, lin2024instance, iclr2025spherepose, cvpr2025spotpose] for estimating object poses within predefined categories without CAD models. These methods fall into two categories: shape prior-based and shape prior-free. Shape prior-based approaches [wang2019normalized, tian2020shape, lin2022category] utilize CAD-derived priors for alignment or direct regression of poses. In contrast, shape prior-free methods [lin2023vi, lin2024instance, chen2024secondpose, iclr2025spherepose, cvpr2025spotpose] remove reliance on priors, learning features directly from input data to estimate poses. Although category-level methods improve generalization over instance-specific methods, they still struggle to generalize to unseen categories.
@@ 2.3 Relative Object Pose Estimation
To enhance the generalization of pose estimation, recent work [zhang2022relpose, zhao20233d, lin2024relpose++, zhao2024dvmnet] has focused on relative pose estimation for unseen objects. Unlike instance-specific methods, these approaches require only a single reference image of a novel object to estimate the relative pose of image, making them highly suitable for applications where data acquisition is costly. Current methods can be broadly categorized into three types: 2D correspondence-based methods [sun2021loftr, sarlin2020superglue, lowe2004distinctive], hypothesis-and-verification-based methods [zhang2022relpose, lin2024relpose++, zhao20233d, nguyen2024nope], and 3D correspondence-based methods [zhao20233d, zhao2024dvmnet]. 2D correspondence-based methods [sun2021loftr, sarlin2020superglue, lowe2004distinctive] establish correspondences between keypoints from both images and use these to compute relative pose. Learned feature-based approaches like SuperGlue [sarlin2020superglue] and LoFTR [sun2021loftr] have achieved robustness in feature matching under moderate viewpoint changes and lighting variations. However, with only a single reference image and substantial viewpoint differences, these methods struggle due to their sensitivity, significantly affecting accuracy [zhang2022relpose, zhao20233d]. Hypothesis-and-verification-based methods [zhang2022relpose, lin2024relpose++, zhao20233d, nguyen2024nope] mitigate these challenges by generating multiple pose hypotheses over the rotation space and evaluating them through similarity networks, as in RelPose [zhang2022relpose] and RelPose++ [lin2024relpose++]. While effective in handling larger viewpoint differences, these methods require extensive sampling and verification, resulting in high computational costs that restrict their suitability for real-time applications. As an alternative, 3D correspondence-based methods [zhao2024dvmnet, zhao20233d] lift 2D features into 3D voxel features and then conduct 3D-3D matching to estimate the pose. For example, DVMNet [zhao2024dvmnet] uses lifted 3D voxel features to facilitate matching even in invisible regions. However, it’s difficult to infer 3D features solely from 2D surface image features, which leads to unreliable 3D matching. Moreover, the dense 3D-3D matching incurs high computational costs due to the cubic complexity.
@@ 3.1 Overview
We tackle the problem of estimating the relative pose between a query image and a reference image belonging to previously unseen object categories. Specifically, we denote the set of object categories available during training as $\omega_{\text{train}}$ and the set of categories used during testing as $\omega_{\text{test}}$, where $\omega_{\text{train}}\cap\omega_{\text{test}}=\emptyset$. This setup introduces several key challenges, including significant viewpoint variations between the images, minimal overlapping regions, and the generalization to novel object categories. Since the translation component of 6D pose can be reliably estimated through existing 2D detection techniques [he2017mask, kirillov2023segment], we focus on addressing the more challenging 3D rotation estimation, which is the same with the mainstream methods [zhang2022relpose, zhao20233d, zhao2024dvmnet].
The framework of our method is shown in Figure 2 (a). Given the query image $I_{q}$ and the reference image $I_{r}$, we first employ a shared feature extractor to obtain feature maps $\mathbf{F}_{q}$ and $\mathbf{F}_{r}$. These feature maps are subsequently passed through symmetric attention blocks [vaswani2017attention] to enhance the image features mutually. Afterward, we use the proposed structure-aware keypoint extraction module to extract keypoints independently from each image. Then, by leveraging our proposed structure-aware correspondence estimation module, we facilitate both intra-image and inter-image feature interactions for the keypoints extracted from the query image, enabling structure-aware feature aggregation. With these updated keypoint features, we lift 2D keypoints to 3D space and regress their corresponding 3D coordinates within the reference coordinate system. Finally, based on the established correspondences, the relative rotation $\Delta\mathbf{R}$ is obtained via a weighted Singular Value Decomposition (wSVD) algorithm [besl1992method].
@@ 3.2 Feature Extraction
We first utilize a pre-trained backbone [weinzaepfel2023croco] to extract image features from both the query and reference images, yielding feature maps $\mathbf{F}_{q},\mathbf{F}_{r}\in\mathbb{R}^{H\times W\times C}$. These feature maps are then processed through a two-step attention mechanism with Multi-Head Self-Attention (MHSA) and Multi-Head Cross-Attention (MHCA) layers to update features:
@@
The same operations are performed for $\mathbf{F}_{r}$. Unlike previous works [zhao20233d, zhao2024dvmnet], the parameters in the MHSA and MHCA modules are shared across the query ($q$) and reference ($r$) images, ensuring consistency between the learned features. After $L$ layers of such interactions, we obtain the final feature maps $\mathbf{F}_{q}^{(L)}$ and $\mathbf{F}_{r}^{(L)}$.
@@ 3.3 Structure-Aware Keypoint Extraction
where $\phi_{l}(\cdot)$ denotes the feature map from the $l$-th layer of the VGG network [simonyan2014very], and $\lambda_{1},\lambda_{2}$ are weighting factors. The $L_{2}$ loss ensures pixel accuracy, while the perceptual loss maintains semantic consistency. By reconstructing the image, we can optimize keypoint distribution end-to-end, ensuring these keypoints cover semantically rich regions of the object’s surface and enhance structural representation.
@@ 3.5 Pose Estimation via 3D Correspondences
After obtaining the 3D keypoint correspondences, following previous work [zhao2024dvmnet], we employ a weighted Singular Value Decomposition (wSVD) approach to solve the relative rotation $\Delta\textbf{R}$:
@@
To solve the minimization problem for estimating the optimal rotation matrix $\Delta\textbf{R}$, we follow the approach of utilizing SVD method to derive the optimal alignment between two sets of 3D keypoints, as used in prior works on point cloud registration [arun1987least]. Specifically, we first compute the covariance matrix $\mathbf{H}$ using the query and reference 3D keypoints along with their confidence scores:
@@
where $q(\Delta\mathbf{R})$ and $q(\Delta\mathbf{R}_{\text{gt}})$ denote the 6D representation for $\Delta\mathbf{R}$ and $\Delta\mathbf{R}_{\text{gt}}$ introduced in  [zhou2019continuity].
@@ 4 Experiment
SuperGlue [sarlin2020superglue] 2D 67.2 45.2 37.7 102.4 15.1 12.1 64.8 26.2 14.3
LoFTR [sun2021loftr] 2D 77.5 37.9 33.1 134.1 9.6 7.7 84.5 24.2 13.5
ZSP [goodwin2022zero] 2D 87.5 25.7 14.6 107.2 4.2 1.5 78.6 10.7 2.7
RelPose [zhang2022relpose] H&V 50.0 64.2 48.6 80.4 20.8 6.7 58.3 26.1 7.0
RelPose++ [lin2024relpose++] H&V 38.5 77.0 69.8 33.5 72.3 42.9 46.6 42.5 15.1
3DAHV [zhao20233d] H&V 3D 28.5 83.5 71.0 28.1 78.6 58.4 41.7 61.5 29.9
DVMNet [zhao2024dvmnet] 3D 19.9 85.9 62.3 20.2 81.5 57.2 36.8 55.1 23.8
@@
Datasets. Following previous works [zhao20233d, zhao2024dvmnet], we evaluate our method on CO3D [reizenstein2021common], Objaverse [deitke2023objaverse] and LineMOD [hinterstoisser2013model], which are widely-used datasets for relative pose estimation. These datasets include diverse synthetic and real data across various object categories. The CO3D dataset contains 18,619 video sequences spanning 51 categories. Following [reizenstein2021common], we train on 41 categories and test on 10 unseen categories to evaluate generalization. The Objaverse dataset consists of synthetic images rendered from 3D models across diverse viewpoints. We select 128 objects for testing and reserve the remaining for training. For LineMOD, we use calibrated real images of 13 household objects. The test set includes 5 objects, which are excluded from training to ensure complete separation.
Implementation Details. The Adam optimizer [kingma2014adam] is employed with an initial learning rate of $2\times 10^{-4}$, which decays by a factor of 0.1 every 200 epochs. The model is trained for 400 epochs with a batch size of 80. All experiments are conducted on 4 NVIDIA RTX3090 GPUs, taking roughly 36 hours. Following previous works [zhang2022relpose, lin2024relpose++, zhao20233d, zhao2024dvmnet], we crop the object from the image by utilizing the ground truth bounding box. Further details can be found in the supplementary material.
Evaluation Metrics. Following previous works [zhao20233d, zhao2024dvmnet], we evaluate our model using two metrics: mean angular error (mAE) and accuracy under predefined thresholds. The angular error $\theta$ between the predicted rotation $\Delta R$ and the ground truth $\Delta R_{\text{gt}}$ is calculated as: